OpenStudy (anonymous):

Solve the equation: 4x2 = 36 A. {6, -6} B. {3, -3} C. {2/3, -2/3} D. {3/2, -3/2}

OpenStudy (lalaly):

4x2-36=0 (2x-6)(2x+6)=0 x=3, x=-3

OpenStudy (anonymous):

B

OpenStudy (lalaly):

its B

OpenStudy (anonymous):

it's b